Evaluate the determinant. - 4 - 3 4 5 6 -1 -2 6 4 0 0 2 4 5 7 1
Solve the equation for x. | 2x - 5 3 2 4 0 -5 5 = 90 The solution set is { }
Answer #1

011 111 A-1666 | 1-3 4-2 -61111 1-3 4-2 R, and Re am some - 1 Al=0© 1-4 -3 4 5 | 6 -1 -2 6 | o o 2 | 1 5 7 1 . Expanding by R3 u1-3 4 51 1-4 -3 41 -1 -2 6-2 6-1 - 21 5 7 1 1 4 5 7 Surt 2017 (3-12X(-44)-16*(-3))+20 X 3 +8X3 – 6X50 - 8X364 - 520 + 496 +60+ 24 -300 - 272 536 ow 99*$ 317-/3 3/4-9% -|- 9 2 [2x5=4*(-5)-*